to buy both the green and blue bikes would cost$1500. to buy the green and red bikes would cost $750. to buy all 3 bikes would cost$2000. how much does each bike cost?
Answer by stanbon(75887)

to buy both the green and blue bikes would cost$1500. to buy the green and red bikes would cost $750. to buy all 3 bikes would cost$2000. how much does each bike cost?
------------
Equations:
g + b + 0 = 1500
g + 0 + r = 750
g + b + r = 2000
---------------------
g = 250
b = 1250
r = 500
=============
Cheers,
Stan H.